OpenStudy (anonymous):

find an equation of the line having the given slope and containg the given point m=2/3,(4,-2)

OpenStudy (anonymous):

Do you know if the question wants point slope or slope intercept form?

OpenStudy (anonymous):

using (y-yo)=m(x- xo) you will get a answer..

OpenStudy (anonymous):

slope intercept form

OpenStudy (anonymous):

y=(2/3)x-14/3

OpenStudy (anonymous):

There are a few ways to do it, probably the easiest is just to plug m,x,y into the equation y=mx+b and solve for b.

OpenStudy (anonymous):

when x=0, y=-14/3

OpenStudy (anonymous):

i can't seem to get the last part

OpenStudy (anonymous):

this "when x=0, y=-14/3"?

OpenStudy (anonymous):

Thanks LeopoldBr

OpenStudy (anonymous):

So you need to know m and b to have a slope intercept equation, right?

OpenStudy (anonymous):

and everyone else for the help

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

2x-3y+4=0